two wires are made up of two different materials whose specific resistance are in the ratio 2:3,lenths are in ratio of 3:4 and area of cross-sections are in the ratio 4:3.find the ratio of their resistances
1
Expert's answer
2012-06-01T08:44:17-0400
R = p · l / A
R1 : R2 = (p1 / p2) · (l1 / l2) · (A2 / A1) = (2 / 3) · (3 / 4)
· (3 /
4) = 18 / 48 = 3 / 8

Answer: 3:8.